# ProactiveMobile: A Comprehensive Benchmark for Boosting Proactive Intelligence on Mobile Devices

### v1

@@ 1 Introduction
Figure 1: A comparison of proactive and reactive paradigm in mobile agents.
@@
However, these agents share a fundamental constraint: they are confined to a reactive paradigm, functioning as passive executors of direct user commands [34, 10]. These models places the entire cognitive burden on the user, from need identification to goal articulation [30], thereby relegating the agent to the role of a high-level tool and fundamentally limiting its potential for seamless integration into daily life [23].
@@
To establish baselines on our benchmark, we fine-tuned Qwen2.5-VL-7B-Instruct [3] and MiMo-VL-7B-SFT-2508 [35] on the training set. We then evaluated its performance on ProactiveMobile alongside a suite of leading closed-source models, including o1 [17], GPT-5 [29], and Gemini-2.5-Pro [7]. Our fine-tuned Qwen2.5-VL-7B-Instruct achieves a success rate of 19.15% on the exact function sequence matching, significantly outperforming closed‑source models, including o1 (15.71%), GPT-5 (7.39%), and Gemini-2.5-Pro (8.91%).
@@ 2.1 LLM-Based Mobile Interaction
Another critical area is task planning and execution. LLMs excel at decomposing high-level natural language commands into a series of executable actions. However, methods based on static prompting often struggled with long-horizon tasks and dynamic environments [51, 34, 42, 40]. Some researches explore fine-tuning or reinforcement learning to enhance the reasoning and prediction capabilities of MLLMs in relative tasks [39, 53, 28, 27, 13]. The maturation of the field is also marked by comprehensive benchmarks [25, 50, 54, 55, 53], which provide standardized environments for evaluating agent performance on realistic tasks.
@@ 3.1 Task Definition
We define proactive intelligence as the task of predicting user’s latent intentions based on their user profile, device status, world information, and behavioral trajectories. The details of these four categories are as follows:
@@ 3.2.1 Acquisition of Behavioral Trajectories
AITZ [50] Largest dataset in android GUI navigation field 9,413
@@
Table 1: Summary of GUI dataset.
@@ 4.5 Ablation Study
To validate our training and output format (Recommendation + Function), we conducted an ablation study comparing it with variants that either omitted the recommendation or added an explicit Think step. The results in Table 4 are decisive. Our chosen strategy achieves the highest SR (19.15%), indicating that compelling the model to articulate a user-facing intent acts as an effective reasoning scaffold. Critically, formats trained without this textual recommendation (Function only and Think + Function) exhibit a catastrophic failure in safety, with False Prediction (FTR) rates near 100%. This demonstrates that generating the intent is indispensable for teaching the model the crucial skill of when not to act.
The study also reveals a crucial trade-off between SR and safety. While adding a Think step (Think + Recommendation + Function) slightly lowered SR, it drastically enhanced safety, slashing the FTR rate to a remarkable 2.21%. This highlights the Think step as a promising direction for building maximally safe agents. Nevertheless, our primary Recommendation + Function approach offers the best-performing balance between SR and reliability, thus validating our core design choice. Further ablation studies, including an analysis of the impact of different contextual dimensions, are detailed in Appendix.
@@ 5 Conclusion
In this work, we address the critical bottleneck hindering the transition of mobile agents from a reactive to a proactive paradigm: the lack of a executable, objective, and realistic benchmark. We introduce ProactiveMobile, a comprehensive benchmark that formalizes the proactive task around a four-dimensional context model, incorporates multi-answer annotations, and uniquely mandates an executable function-call sequence output. Our extensive experiments validate that proactivity is a specialized, learnable capability. This is consistently demonstrated as fine-tuning on our benchmark boosts different models’ performance, with our top-performing model achieving a 19.15% success rate—establishing a new state-of-the-art that surpasses even leading proprietary models like o1 (15.71%). This proves the efficacy of ProactiveMobile as an essential tool for targeted training and highlights the significant gap in current models’ out-of-the-box abilities.

### v2

@@ 1 Introduction
Figure 1: A comparison of proactive and reactive paradigms in mobile agents.
@@
However, these agents share a fundamental constraint: they are confined to a reactive paradigm, functioning as passive executors of direct user commands [34, 10]. These models place the entire cognitive burden on the user, from need identification to goal articulation [30], thereby relegating the agent to the role of a high-level tool and fundamentally limiting its potential for seamless integration into daily life [23].
@@
To establish baselines on our benchmark, we fine-tuned Qwen2.5-VL-7B-Instruct [3] and MiMo-VL-7B-SFT-2508 [35] on the training set. We then evaluated their performance on ProactiveMobile alongside a suite of leading closed-source models, including o1 [17], GPT-5 [29], and Gemini-2.5-Pro [7]. Our fine-tuned Qwen2.5-VL-7B-Instruct achieves a success rate of 19.15% on the exact function sequence matching, significantly outperforming closed‑source models, including o1 (15.71%), GPT-5 (7.39%), and Gemini-2.5-Pro (8.91%).
@@ 2.1 LLM-Based Mobile Interaction
Another critical area is task planning and execution. LLMs excel at decomposing high-level natural language commands into a series of executable actions. However, methods based on static prompting often struggle with long-horizon tasks and dynamic environments [51, 34, 42, 40]. Some research explores fine-tuning or reinforcement learning to enhance the reasoning and prediction capabilities of MLLMs in related tasks [39, 53, 28, 27, 13]. The maturation of the field is also marked by comprehensive benchmarks [25, 50, 54, 53], which provide standardized environments for evaluating agent performance on realistic tasks.
@@ 3.1 Task Definition
We define proactive intelligence as the task of predicting users’ latent intentions based on their user profile, device status, world information, and behavioral trajectories. The details of these four categories are as follows:
@@ 3.2.1 Acquisition of Behavioral Trajectories
AITZ [50] Largest dataset in the Android GUI navigation field 9,413
@@
Table 1: Summary of GUI datasets.
@@ 4.5 Ablation Study
To validate our training and output format (Recommendation + Function), we conducted an ablation study comparing it with variants that either omitted the recommendation or added an explicit Think step. The results in Table 4 are decisive. Our chosen strategy achieves the highest SR (19.15%), indicating that compelling the model to articulate a user-facing intent acts as an effective reasoning scaffold. Critically, formats trained without this textual recommendation (Function only and Think + Function) exhibit a catastrophic failure in safety, with False Trigger Rate (FTR) rates near 100%. This demonstrates that generating the intent is indispensable for teaching the model the crucial skill of when not to act.
The study also reveals a crucial trade-off between SR and safety. While adding a Think step (Think + Recommendation + Function) slightly lowered SR, it drastically enhanced safety, slashing the FTR rate to a 2.21%. This highlights the Think step as a promising direction for building maximally safe agents. Nevertheless, our primary Recommendation + Function approach offers the best-performing balance between SR and reliability, thus validating our core design choice. Further ablation studies, including an analysis of the impact of different contextual dimensions, are detailed in Appendix.
@@ 5 Conclusion
In this work, we address the critical bottleneck hindering the transition of mobile agents from a reactive to a proactive paradigm: the lack of an executable, objective, and realistic benchmark. We introduce ProactiveMobile, a comprehensive benchmark that formalizes the proactive task around a four-dimensional context model, incorporates multi-answer annotations, and uniquely mandates an executable function-call sequence output. Our extensive experiments validate that proactivity is a specialized, learnable capability. This is consistently demonstrated as fine-tuning on our benchmark boosts different models’ performance, with our top-performing model achieving a 19.15% success rate—establishing a new state-of-the-art that surpasses even leading proprietary models like o1 (15.71%). This demonstrates the efficacy of ProactiveMobile as an essential tool for targeted training and highlights the significant gap in current models’ out-of-the-box abilities.